# ProtChatGPT: Towards Understanding Proteins with Large Language Models

### v1

@@ Appendix C Details on Metrics
In our implementation, we use seven different metrics on the proposed ProtChatGPT to verify the performance. These metrics not only focus on the low-level lexical and syntactic alignment but also on high-level semantic information.
@@
$p_{n}=\frac{\sum\limits_{C\in\{Candidates\}}\sum\limits_{\textit{n-gram}\in C}Count_{clip}(\textit{n-gram})}{\sum\limits_{C^{\prime}\in\{Candidates\}}\sum\limits_{\textit{n-gram}^{\prime}\in C^{\prime}}Count(\textit{n-gram}^{\prime})}$ (6)
$\text{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases}$ (7)
$\text{BLEU}=\text{BP}\cdot exp\left(\sum_{n=1}^{N}w_{n}logp_{n}\right)$ (8)
@@
ROUGE-L (Lin & Hovy, 2002) finds frequent use in the automatic evaluation of text summarization and machine translation. It calculates the longest common subsequence between the candidate and reference texts, prioritizing recall over precision. ROUGE-L can be expressed mathematically as:
$R_{lcs}=\frac{LCS(X,Y)}{m}$ (9)
$P_{lcs}=\frac{LCS(X,Y)}{n}$ (10)
@@
In this context, $X$ represents the predicted text with a length of $n$. $Y$ represents the ground truth text with a length of $m$. $\beta$ is a hyperparameter used to adjust the emphasis on precision and recall. LCS calculates the length of the longest common subsequence, $R_{lcs}$ measures recall, and $P_{lcs}$ measures precision, respectively.
METEOR (Banerjee & Lavie, 2005) provides a comprehensive evaluation of machine-generated text by considering not only exact word matches but also synonyms and stemming. It combines precision, recall, and alignment factors to offer a holistic assessment. METEOR is mathematically represented as:
$F=\frac{(\alpha^{2}+1)P}{R+\alpha P}$ (12)
$\text{Meteor}=(1-Penalty)\cdot F$ (13)
@@
CIDEr (Vedantam et al., 2015) (Consensus-based Image Description Evaluation) primarily assesses the quality of image captions produced by automated systems. It places importance on consensus among multiple reference captions and emphasizes the inclusion of diverse descriptive words.
The mathematical formulation of CIDEr is given by:
$CIDEr_{n}(c,S)=\frac{1}{M}\sum_{i=1}^{M}\frac{g^{n}(c)\cdot g^{n}(S_{i})}{||g^{n}(c)||\times||g^{n}(S_{i})||}$ (14)
In the formula, $c$ represents the candidate text, $S$ denotes the set of reference texts, $n$ specifies the use of n-grams, $M$ represents the number of reference texts, and $g$ corresponds to the TF-IDF vector based on n-grams.
SPICE (Anderson et al., 2016) offers a metric designed to evaluate the semantic content of image captions, with a focus on their precision within generated captions.
SPICE’s mathematical expression is as follows:
$P(c,S)=\frac{|T(G(c))\otimes T(G(S))|}{|T(G(c))|}$ (15)
$R(c,S)=\frac{|T(G(c))\otimes T(G(S))|}{|T(G(S))|}$ (16)
$SPICE(c,S)=\frac{2\cdot P(c,S)\cdot R(c,S)}{P(c,S)+R(c,S)}$ (17)

### v2

@@ Appendix I Details on Metrics
In our implementation, we use seven different metrics on the proposed ProtChatGPT to verify the performance. These metrics focus not only on low-level lexical and syntactic alignment but also on high-level semantic information.
@@
(5) $p_{n}=\frac{\sum\limits_{C\in\{Candidates\}}\sum\limits_{\textit{n-gram}\in C}Count_{clip}(\textit{n-gram})}{\sum\limits_{C^{\prime}\in\{Candidates\}}\sum\limits_{\textit{n-gram}^{\prime}\in C^{\prime}}Count(\textit{n-gram}^{\prime})},$
(6) $\text{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r,\end{cases}$
(7) $\text{BLEU}=\text{BP}\cdot exp\left(\sum_{n=1}^{N}w_{n}logp_{n}\right),$
@@
ROUGE-L (Lin and Hovy, 2002) finds frequent use in the automatic evaluation of text summarization and machine translation. It calculates the longest common subsequence between the candidate and reference texts, prioritizing recall over precision. ROUGE-L can be formulated as:
(8) $R_{lcs}=\frac{LCS(X,Y)}{m},$
(9) $P_{lcs}=\frac{LCS(X,Y)}{n},$
@@
where $X$ represents the predicted text with a length of $n$. $Y$ represents the ground truth text with a length of $m$. $\beta$ is a hyperparameter used to adjust the emphasis on precision and recall. LCS calculates the length of the longest common subsequence, $R_{lcs}$ measures recall, and $P_{lcs}$ measures precision, respectively.
METEOR (Banerjee and Lavie, 2005) provides a comprehensive evaluation of machine-generated text by considering not only exact word matches but also synonyms and stemming. It combines precision, recall, and alignment factors to offer a holistic assessment. METEOR is formulated as:
(11) $F=\frac{(\alpha^{2}+1)P}{R+\alpha P},$
(12) $\text{Meteor}=(1-Penalty)\cdot F,$
@@
CIDEr (Vedantam et al., 2015) (Consensus-based Image Description Evaluation) primarily assesses the quality of image captions produced by automated systems. It places importance on consensus among multiple reference captions and emphasizes the inclusion of diverse descriptive words. The mathematical formulation of CIDEr is given by:
(13) $CIDEr_{n}(c,S)=\frac{1}{M}\sum_{i=1}^{M}\frac{g^{n}(c)\cdot g^{n}(S_{i})}{||g^{n}(c)||\times||g^{n}(S_{i})||},$
where $c$ represents the candidate text, $S$ denotes the set of reference texts, $n$ specifies the use of n-grams, $M$ represents the number of reference texts, and $g$ corresponds to the TF-IDF vector based on n-grams.
SPICE (Anderson et al., 2016) offers a metric designed to evaluate the semantic content of image captions, with a focus on their precision within generated captions. The formulation of SPICE is as follows:
(14) $P(c,S)=\frac{|T(G(c))\otimes T(G(S))|}{|T(G(c))|},$
(15) $R(c,S)=\frac{|T(G(c))\otimes T(G(S))|}{|T(G(S))|},$
(16) $SPICE(c,S)=\frac{2\cdot P(c,S)\cdot R(c,S)}{P(c,S)+R(c,S)},$